# Tree-Planted Transformers: Large Language Models with Implicit Syntactic Supervision

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success thanks to scalability on large text corpora, but have some drawback in training efficiency. In contrast, Syntactic Language Models (SLMs) can be trained efficiently to reach relatively high performance thanks to syntactic supervision, but have trouble with scalability. Thus, given these complementary advantages of LLMs and SLMs, it is necessary to develop an architecture that integrates the scalability of LLMs with the training efficiency of SLMs, namely Syntactic Large Language Models (SLLM). In this paper, we propose a novel method dubbed tree-planting: implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language. Specifically, Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding. Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit syntactic supervision, significantly outperformed various SLMs with explicit syntactic supervision that generate hundreds of syntactic structures in parallel, suggesting that tree-planting and TPTs are the promising foundation for SLLMs.
Tree-Planted Transformers:
Large Language Models with Implicit Syntactic Supervision
@@ 1 Introduction
Recent years have witnessed remarkable success in Large Language Models (LLMs) based on Transformer LMs (Vaswani et al., 2017). The success of LLMs suggests that continual learning on large text corpora is essential for LMs to acquire a wide range of world knowledge and solve various downstream tasks. However, despite their success, LLMs have some drawback in training efficiency. For example, GPT-3 (Brown et al., 2020) is trained on around larger data than a 12-year-old human would have experienced (Warstadt et al., 2023), indicating that LLMs require tremendous training corpus and computational resources.
@@
On another strand, previous work has revealed that Syntactic Language Models (SLMs), the integration of LMs with explicit syntactic supervision, can achieve high performance under data-constrained settings (Dyer et al., 2016; Noji and Oseki, 2021; Qian et al., 2021; Sartran et al., 2022; Yoshida and Oseki, 2022; Murty et al., 2023). For example, Sartran et al. (2022) showed that some SLMs can achieve comparable syntactic knowledge to an LLM-like model111Due to the rapid advances in recent years, what were once considered LLMs are no longer deemed “large” by current standards. We will refer to Transformer LMs larger than or equal to GPT-2 (Radford et al., 2018) as LLM-like models. that is trained with medium—around larger—data, suggesting that syntactic supervision is essential for LMs to achieve high training efficiency. However, despite their training efficiency, SLMs have trouble with scalability: small SLMs cannot compete with LLMs trained on larger data. Thus, given these complementary advantages of LLMs and SLMs, it is necessary to develop an architecture that integrates the scalability of LLMs with the training efficiency of SLMs, namely Syntactic Large Language Models (SLLM; Table 1).
@@
In this paper, we propose a novel method dubbed tree-planting:222The term “tree-planting” coincidentally bears a resemblance to the term used in Mueller and Linzen (2023), but this work diverges from ours in its motivation. Specifically, Mueller and Linzen (2023) investigated biases that enable syntactic generalization in Transformer LMs, from the perspectives of architectural features (depth, width, and number of parameters), as well as the genre and size of training corpus. implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language. Specifically, Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding (Figure 1). Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit syntactic supervision, significantly outperformed various SLMs with explicit syntactic supervision that generate hundreds of syntactic structures in parallel, suggesting that tree-planting and TPTs are the promising foundation for SLLMs.
2 Background
2.1 Large Language Model
Large Language Models are typically based on Transformer LMs (Vaswani et al., 2017) with a large number of parameters and trained on vast amounts of data. A major reason that Transformer LMs are employed as the base architecture for LLMs is their self-attention mechanism, which enables efficient parallel computation on GPUs.
The self-attention mechanism of Transformer LMs computes a representation for predicting the next token through a weighted sum of each token in the context. Specifically, when predicting the -th token, the attention weights from the -th token to the -th token is computed as follows:
where and represent the query vector of the -th token and the key vector of the -th token, respectively, and denotes the dimension of the key vector. As Equation 1 shows, the computation for the -th token prediction does not depend on any computation for the -th token predictions, which enables efficient parallel computation. This property of the self-attention mechanism enables the development of LLMs but it is important to note that these models do not employ any syntactic supervision, although syntactic structures are one of the fundamental properties of natural languages Chomsky (1957).
| Scalability | Training efficiency | |
| LLM | ✓ | |
| SLM | ✓ | |
| SLLM | ✓ | ✓ |
@@ 2.2 Syntactic Language Model
Syntactic Language Models (SLMs) are a generative model of a token sequence and its syntactic structure :
where denotes the sequence of actions to generate both the token sequence and syntactic structure. For example, in top-down and left-to-right SLMs, each could be either generating a token, opening a constituent, or closing a constituent.
Recently, several SLMs based on the Transformer architecture have been proposed, achieving higher syntactic knowledge than medium LLM-like models Qian et al. (2021); Sartran et al. (2022); Murty et al. (2023). However, because SLMs model the joint probability of a token sequence and its syntactic structure, they cannot be trained on other than treebanks, which prevents them from scaling on large text corpora Moreover, SLMs also have practical drawback in inference costs: when utilized as LMs, they require hundreds of syntactic structures via beam search (Stern et al., 2017; Crabbé et al., 2019) or an external parser, to marginalize joint distribution to precisely approximate .
@@ 2.3 Constraints on attention weights
As discussed in Subsection 2.2, the bottleneck that prevents SLMs from scaling on large text corpora is their modeling space of the joint probability. To achieve the foundational architecture for SLLMs, it is necessary to introduce syntactic knowledge without changing the modeling space of their underlying Transformer LMs. For our goal, we will build upon another line of approach that constrains attention weights according to syntactic structures—typically targeting bidirectional Transformer Encoders like BERT (Devlin et al., 2019)—and extend it to unidirectional Transformer LMs. Table 2 summarizes the previous work in this line of approach, comparing our tree-planting (Section 3) against others based on the requirements for SLLM: (i) parser-free inference, (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation.
@@
First, the majority of these approaches are purely motivated to explicitly restrict attention weights with syntactic structures from external parsers, under the assumption that these parsers would be available during inference (Wu et al., 2018; Nguyen et al., 2020; Bugliarello and Okazaki, 2020; Bai et al., 2021; Sachan et al., 2021; Slobodkin et al., 2022). These studies achieved successful performance in their respective downstream tasks, but not only are they all not directly applicable to unidirectional LMs, they also require external parsers during inference, rendering them unsuitable as the foundation for SLLM.
@@
Second, several studies proposed approaches that do not require external parsers during inference. Wang et al. (2019) aimed at an unsupervised approach, where a hierarchical architectural bias widens the range of neighboring tokens eligible to attend from lower to upper layers, though this method is also not aligned with our goal of achieving higher training efficiency via syntactic supervision. Furthermore, Strubell et al. (2018) and Chen et al. (2023) designed the loss functions that implicitly encourage the attention to syntactic parents or children for each token, satisfying the 3/4 requirements for SLLM. However, this approach is potentially not suitable for unidirectional LMs where the existence of the dependent in the left context is not guaranteed.
@@
Finally, the approach also most closely aligned with the spirit of this research is a hybrid Parser and neural Language Model (PaLM; Peng et al., 2019). PaLM is the integration of an RNN LM with an additional attention layer, which would be supervised to attend the constituent spans among the spans ending at time : . Although PaLM also meets the 3/4 requirements, it was by nature proposed for RNN LMs. The challenge arises when adapting PaLM to Transformer LMs; the generation of embeddings for the spans introduces a significant bottleneck in parallel computation with the self-attention mechanism.
@@
To sum up, none of the previous approaches fully satisfy the requirements for SLLM, highlighting the necessity for innovative methodologies.
@@ 3 Proposed method: tree-planting
In this paper, we propose a novel method dubbed tree-planting: implicitly “plant” trees into attention weights of Transformer LMs to reflect syntactic structures of natural language (Figure 2). Specifically, Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which learn syntax on small treebanks via tree-planting and then scale on large text corpora via continual learning with syntactic scaffolding. Tree-planting is strictly designed to satisfy the requirements for SLLM: (i) parser-free inference, (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation.
@@ 3.1 Supervision of attention weights
In producing the supervision of attention weights, we extend the notion of syntactic distance (Shen et al., 2018, 2019; Du et al., 2020), a 1D sequence of the number of edges on syntactic structures between two consecutive words, to a 2D matrix between all pairs of words:
where and represent the -th and -th words, respectively, and is the function that maps a pair of words to the number of edges on syntactic structures between them. This notion of syntactic distance matrix is theory-neutral: applied to any kind of syntactic structure, as long as the number of edges can be counted on it.333When applied to dependency structures, we ignore the direction of syntactic dependency.
@@
where represents the supervision of the attention weight from the -th word to the -th word when predicting the -th word. This design of the supervision expects the attention weight of each word to decrease exponentially with its number of edges between the predicted word;444We adopt an exponential function as Lin and Tegmark (2017) reported that the mutual information between words will decay exponentially with respect to the number of edges on the syntactic structure between them. this alone successfully satisfies the 3/4 requirements for SLLM: (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation. To fulfill the remaining requirement of (i) parser-free inference, we adopt a strategy similar to that of Strubell et al. (2018); Chen et al. (2023), designing the loss function to implicitly supervise attention.
@@ 3.2 Loss function
The supervision in Subsection 3.1 is produced at the word level but LLMs typically take their input at the subword level. To bridge this gap, we first convert the subword-level attention weight matrix from a targeted Transformer LM to the word-level attention weight matrix as follows:
where represents the word-level attention weight from the -th word to the -th word. is defined as the sum of the subword-level attention weights over the subword inside when predicting the subword inside , with representing the subword-level attention weight from the -th subword to the -th subword and and being the functions that map words to their start and end subword index, respectively. We employ from specific attention heads called tree-planted heads.555Qian et al. (2021) also proposed the architecture which constrains some attention heads based on syntactic structures, or PLM-mask. PLM-mask and our tree-planting are similar in spirit, but they are quite different in their implementation: PLM-mask is a type of SLM that jointly generates a word sequence and its syntactic structure, but tree-planting builds TPTs, a type of LM. Furthermore, PLM-mask explicitly masks the attention weights based on the local parser state but tree-planting implicitly guides attention weights to reflect the whole syntactic structure.
To implicitly supervise the word-level attention weight matrix with the supervision , we introduce a tree-planting loss employing a Kullback–Leibler (KL) Divergence loss :666This loss function is inspired by Ma et al. (2023), which guides attention weights to focus on relevant texts in a document-level relation extraction task.
@@
where represents a tree-planting loss for each tree-planted head , is the total number of tree-planted heads, and is a weight that balances the importance of the next word prediction loss and the average tree-planting loss. Transformer LMs trained with this loss function will be called Tree-Planted Transformers (TPT).
@@ 4 Experiment
To confirm that syntactic knowledge is introduced to TPTs by tree-planting, we conduct training on a small treebank and targeted syntactic evaluations on a syntactic knowledge benchmark.

### v2

# Tree-Planted Transformers: Unidirectional Transformer Language Models with Implicit Syntactic Supervision

@@ Abstract
Syntactic Language Models (SLMs) can be trained efficiently to reach relatively high performance; however, they have trouble with inference efficiency due to the explicit generation of syntactic structures. In this paper, we propose a new method dubbed tree-planting: instead of explicitly generating syntactic structures, we “plant” trees into attention weights of unidirectional Transformer LMs to implicitly reflect syntactic structures of natural language. Specifically, unidirectional Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which inherit the training efficiency from SLMs without changing the inference efficiency of their underlying Transformer LMs. Targeted syntactic evaluations on the SyntaxGym benchmark demonstrated that TPTs, despite the lack of explicit generation of syntactic structures, significantly outperformed not only vanilla Transformer LMs but also various SLMs that generate hundreds of syntactic structures in parallel. This result suggests that TPTs can learn human-like syntactic knowledge as data-efficiently as SLMs while maintaining the modeling space of Transformer LMs unchanged.
[…]
Unidirectional Transformer Language Models
with Implicit Syntactic Supervision
@@ 1 Introduction
Recent years have witnessed remarkable success in Large Language Models (LLMs) based on Transformer LMs (Vaswani et al., 2017). However, despite their success, Transformer LMs have some drawback in training efficiency—especially when compared with humans. For example, GPT-3 (Brown et al., 2020) is trained on around larger data than a 12-year-old human would have experienced (Warstadt et al., 2023), indicating that Transformer LMs lack sufficient inductive bias for language acquisition.
@@
On another strand, previous work has revealed that Syntactic Language Models (SLMs), defined as a generative models of a token sequence and its syntactic structures, can achieve high syntactic performance under data-constrained settings (Dyer et al., 2016; Noji and Oseki, 2021; Qian et al., 2021; Sartran et al., 2022; Yoshida and Oseki, 2022; Murty et al., 2023). For example, Sartran et al. (2022) showed that some SLMs can achieve comparable syntactic performance to an LLM-like model111Due to the rapid advances in recent years, what were once considered LLMs are no longer deemed “large” by current standards. We will refer to Transformer LMs larger than or equal to GPT-2 (Radford et al., 2018) as LLM-like models. that is trained with medium—around larger—data, suggesting that syntactic supervision is essential for LMs to achieve high training efficiency. However, despite their training efficiency, SLMs have trouble with inference efficiency—they require hundreds of syntactic structures generated via beam search (Stern et al., 2017; Crabbé et al., 2019) or an external parser to precisely approximate marginal distribution over a token sequence, which naturally incurs the costs hundreds of times higher than their underlying sequential models.
@@
In this paper, we propose a new method dubbed tree-planting: instead of explicitly generating syntactic structures, we “plant” trees into attention weights of unidirectional Transformer LMs to implicitly reflect syntactic structures of natural language.222The term “tree-planting” coincidentally bears a resemblance to the term used in Mueller and Linzen (2023), but this work diverges from ours in its motivation. Specifically, Mueller and Linzen (2023) investigated biases that enable syntactic generalization in Transformer LMs, from the perspectives of architectural features (depth, width, and number of parameters), as well as the genre and size of training corpus. Specifically, unidirectional Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which inherit the training efficiency from SLMs without changing the inference efficiency of their underlying Transformer LMs (Figure 1).
Previous studies have also explored the syntactic supervision of attention weights, mainly targeting bidirectional Transformer Encoders (Wu et al., 2018; Nguyen et al., 2020; Bugliarello and Okazaki, 2020; Bai et al., 2021; Sachan et al., 2021; Slobodkin et al., 2022; inter alia) (§2). These encoder-oriented approaches assume the entire sentence as input and typically aim to reflect the syntactic relationship between input words in a bottom-up manner. In contrast, tree-planting is uniquely designed for unidirectional Transformer LMs that can be used for text generation—and they have recently increased prevalence because of their compatibility with instruct tuning (Zhang et al., 2024) or Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022). Specifically, tree-planting considers syntactic structures involving the next word to be generated, by focusing on the syntactic distance (Shen et al., 2018, 2019; Du et al., 2020) between the next word and the previous words in the context (§3).
Targeted syntactic evaluations on the SyntaxGym benchmark (Gauthier et al., 2020) demonstrated that TPTs, despite the lack of explicit generation of syntactic structures, significantly outperformed not only vanilla Transformer LMs but also various SLMs that generate hundreds of syntactic structures in parallel. This result suggests that TPTs can learn human-like syntactic knowledge as data-efficiently as SLMs while maintaining the modeling space of Transformer LMs unchanged. Furthermore, closer inspection of syntactic phenomena implied that tree-planting shows high compatibility with dependency structures (§4).
Additionally, we analyzed the two hyperparameters introduced by tree-planting: (i) the number of heads where the tree-planting loss is applied, and (ii) the balance between the next-word prediction loss and the tree-planting loss. Our results demonstrated that the highest accuracy was generally achieved when a single head was adopted as a tree-planted head while the balancing parameter emerged as a crucial hyperparameter; excessively high and low weights on the tree-planting loss both led to ineffective outcomes (§5).
2 Related work
@@ 2.1 Syntactic Language Model
Syntactic Language Models (SLMs) are a generative model of a token sequence and its syntactic structure . Formally, SLMs are defined as:
where denotes the sequence of actions to generate both the token sequence and syntactic structure. For example, in top-down and left-to-right SLMs, each could be either generating a token or opening/closing a constituent.
Recently, several SLMs based on the Transformer architecture have been proposed, achieving higher syntactic performance than medium LLM-like models Qian et al. (2021); Sartran et al. (2022); Murty et al. (2023). However, because SLMs generate both a syntactic structure and token sequence, they cannot be directly utilized as LMs, or a generative model of a token sequence. To precisely approximate the marginal distribution over a token sequence, i.e., (where represents the set of possible syntactic structures behind ), they require hundreds of syntactic structures generated via beam search (Stern et al., 2017; Crabbé et al., 2019) or an external parser. Although actual costs would depend on the hardware and metrics, the calculation of for each structure in naturally multiplies inference costs by compared to the sequential models that directly calculates . Furthermore, additional costs would be incurred by the beam search procedure or the external parser itself.
@@ 2.2 Constraints on attention weights
As discussed in §2.1, the bottleneck that impairs SLMs’ inference efficiency is their modeling space of the joint probability. To achieve the architecture that inherits the training efficiency of SLMs without changing the inference efficiency of their underlying Transformer LMs, it is necessary to introduce syntactic supervision without changing the modeling space of the sequential models. For our goal, we will build upon another line of approach that constrains attention weights according to syntactic structures—mainly targeting bidirectional Transformer Encoders like BERT (Devlin et al., 2019).333Beyond studies that constrain attention weights according to syntactic structures, there are also investigations that aimed at incorporating various types of information, such as word alignment, into the attention mechanisms (e.g., Yin et al., 2021). This subsection, however, specifically focuses on studies that target syntactic biases. Table 1 summarizes the previous work in this line of approach, comparing our tree-planting (§3) against others based on the requirements for our goal: (i) parser-free inference, (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation.
@@
First, the majority of these approaches are purely motivated to explicitly restrict attention weights with syntactic structures from external parsers, under the assumption that these parsers would be available during inference (Wu et al., 2018; Nguyen et al., 2020; Bugliarello and Okazaki, 2020; Bai et al., 2021; Sachan et al., 2021; Slobodkin et al., 2022). These studies achieved successful performance in their respective downstream tasks, but not only are their approaches all not directly applicable to unidirectional LMs, they also require external parsers during inference, rendering them not aligned with our goal of the inference efficient architecture.
@@
Second, several approaches have been proposed that eliminate the need for external parsers during inference, but they still fall short of meeting all the requirements. Wang et al. (2019) aimed at an unsupervised approach, where a hierarchical architectural bias widens the range of neighboring tokens eligible to attend from lower to upper layers, yet this method is still not aligned with our goal of achieving higher training efficiency via syntactic supervision. Additionally, Strubell et al. (2018) and Chen et al. (2023) designed the loss functions that implicitly encourage the attention to syntactic parents or children for each token, satisfying the 3/4 requirements for our goal.444Other than the approach to constrain attention weights, Tziafas et al. (2023) proposed a method to train Transformer Encoders in a multi-task setting of masked language modeling and categorial grammar supertagging. As an anonymous reviewer correctly pointed out, this method also satisfied the 3/4 requirements: (i), (ii), and (iv). However, these approaches are potentially encoder-oriented and not suitable for unidirectional LMs; they assume the entire sentence as input and reflect the syntactic relationship between input words in a bottom-up manner.
@@
Finally, another approach also closely aligned with the spirit of this research is a hybrid Parser and neural Language Model (PaLM; Peng et al., 2019). PaLM is the integration of an unidirectional RNN LM with an additional attention layer, which would be supervised to attend the constituent spans among the spans ending at time : . Although PaLM also meets the 3/4 requirements, it was by nature proposed for RNN LMs. The challenge arises when adapting PaLM to Transformer LMs; the generation of embeddings for the spans cannot be parallelized in a manner that is compatible with a self-attention mechanism of Transformer LMs.
@@
To sum up, none of the previous approaches fully satisfy the requirements for our goal, highlighting the necessity for new methodologies.
@@ 3 Proposed method: tree-planting
In this paper, we propose a new method dubbed tree-planting: we “plant” trees into attention weights of unidirectional Transformer LMs to reflect syntactic structures of natural language (Figure 2). Specifically, unidirectional Transformer LMs trained with tree-planting will be called Tree-Planted Transformers (TPT), which inherit the training efficiency from SLMs without changing the inference efficiency of their underlying Trans- former LMs. Tree-planting is uniquely designed for unidirectional Transformer LMs, satisfying all the requirements discussed in §2.2: (i) parser-free inference, (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation.
@@ 3.1 Supervision of attention weights
A self-attention mechanism of unidirectional Transformer LMs computes a representation for predicting the next token through a weighted sum of each token in the context. Specifically, when predicting the -th token, the attention weights from the -th token to the -th token is computed as follows:
where and represent the query vector of the -th token and the key vector of the -th token, respectively, and denotes the dimension of the key vector. As Equation 2 shows, the computation for the -th token prediction does not depend on any computation for the -th token predictions, which enables parallel computation.
In producing the supervision of attention weights, we extend the notion of syntactic distance array (Shen et al., 2018, 2019; Du et al., 2020), a 1D array of the number of edges on syntactic structures between two consecutive words, to a 2D matrix between all pairs of words:
where and represent the -th and -th words, respectively, and is the function that maps a pair of words to the number of edges on syntactic structures between them. This notion of syntactic distance matrix could be applied to any kind of syntactic structure, as long as the number of edges can be counted on it.555Note importantly that compatibility with the syntactic distance matrix varies across types of syntactic structures because some kind of information would be lost during the conversion from syntactic structures to the matrix. For example, when applied to dependency structures, the information on the direction of syntactic dependency will be lost.
@@
where represents the supervision of the attention weight from the -th word to the -th word when predicting the -th word. In short, this supervision expects the attention weight of each word to decrease exponentially with the number of edges between the predicted word.666We adopt an exponential function as Lin and Tegmark (2017) reported that the mutual information between words would decay exponentially with respect to the number of edges on the syntactic structure between them.
This supervision design is highly oriented towards unidirectional Transformer LMs; it considers syntactic structures involving the next word to be generated in a manner compatible with parallel computation of the self-attention mechanism. This only successfully satisfies the 3/4 requirements for our purpose: (ii) syntactic supervision, (iii) unidirectional LM, and (iv) parallel computation. To fulfill the remaining requirement of (i) parser-free inference, we adopt a strategy similar to that of Strubell et al. (2018); Chen et al. (2023), designing the loss function to implicitly supervise attention.
@@ 3.2 Loss function
The supervision in §3.1 is produced at the word level but Transformer LMs typically take their input at the subword level. To bridge this gap, we first convert the subword-level attention weight matrix from a targeted Transformer LM to the word-level attention weight matrix as follows:
where represents the word-level attention weight from the -th word to the -th word. is defined as the sum of the subword-level attention weights over the subword inside when predicting the subword inside , with representing the subword-level attention weight from the -th subword to the -th subword and and being the functions that map words to their start and end subword index, respectively. We employ from specific attention heads called tree-planted heads.777Qian et al. (2021) also proposed the architecture which constrains some attention heads based on syntactic structures, or PLM-mask. PLM-mask and our tree-planting are similar in spirit, but they are quite different in their implementation: PLM-mask is a type of SLM that jointly generates a word sequence and its syntactic structure, but tree-planting builds TPTs, a type of LM. Furthermore, PLM-mask explicitly masks the attention weights based on the local parser state but tree-planting implicitly guides attention weights to reflect the whole syntactic structure.
To implicitly supervise the word-level attention weight matrix with the supervision , we introduce a tree-planting loss employing a Kullback–Leibler (KL) Divergence loss :888This loss function is inspired by Ma et al. (2023), which guides attention weights to focus on relevant texts in a document-level relation extraction task.
@@
where represents a tree-planting loss for each tree-planted head , is the total number of tree-planted heads, and is a weight that balances the importance of the next word prediction loss and the average tree-planting loss. Unidirectional Transformer LMs trained with this loss function will be called Tree-Planted Transformers (TPT).
@@ 4 Experiment
To investigate whether TPTs can learn human-like syntactic knowledge as data-efficiently as SLMs while maintaining the modeling space of Transformer LMs unchanged, we conduct training on a small treebank and targeted syntactic evaluations on a syntactic knowledge benchmark.